# Metric-aware LLM inference

### v1

@@ Abstract
Large language models (LLMs) have demonstrated strong results on a range of NLP tasks. Typically, outputs are obtained via autoregressive sampling from the LLM’s underlying distribution. We show that this inference strategy can be suboptimal for a range of tasks and associated evaluation metrics. As a remedy, we propose metric aware LLM inference: a decision theoretic approach optimizing for custom metrics at inference time. We report improvements over baselines on academic benchmarks and publicly available models.
@@ 1 Introduction
Large language models (LLMs) are currently the most capable models across many NLP tasks (OpenAI et al., 2023; Google and et al., 2023; Touvron et al., 2023; Gemini Team and et al., 2023). Owing to their remarkable few- and zero-shot abilities (Wei et al., 2022; Kojima et al., 2023), many tasks can be addressed without conducting any additional model training on in-domain datasets: instead, users can obtain useful results by querying a pretrained LLM with a suitably crafted input prompt.
Different NLP tasks employ different evaluation metrics to assess the quality of an LLM. One popular metric is the exact match (EM), which penalises any response not exactly equal to that provided in the dataset annotation. This is an analogue of the conventional classification accuracy.
While the EM is an intuitive metric, there are many cases where it is not suitable. For example, metrics such as squared error, mean absolute error or correlation scores are often preferred on tasks where outputs are numerical, or where categories are ordered. Such tasks occur in prominent NLP challenges such as relevance scoring (Cer et al., 2017) and sentiment analysis (Fathony et al., 2017). Further, BLEU (Papineni et al., 2002) and ROUGE (Ganesan, 2018) scores are popular for natural language tasks such as summarization and machine translation, while $F_{1}$ score is used to evaluate reading comprehension (Joshi et al., 2017) and text segmentation (Lukasik et al., 2020). Moreover, as noted in Schaeffer et al. (2023), the EM score may not properly reflect improvements in model performance due to its discontinuity.
@@
Despite the wide variety of evaluation metrics, LLM inference is typically performed in the same manner for every task: namely, one performs auto-regressive sampling from the LLM’s underlying distribution (see §2). While intuitive, such inference notably does not explicitly consider the downstream evaluation metric of interest. This raises a natural question: is there value in adapting the inference procedure to the evaluation metric at hand?
In this paper, we focus on the zero-shot abilities of LLMs on tasks evaluated on different metrics. We make a key observation that choosing the most likely target for the output corresponds to inherently optimizing for the exact match evaluation metric; consequently, this inference strategy is not optimal when the EM is not the metric of interest. As a remedy to this, we propose metric aware LLM inference (MALI), a decision-theoretic approach where a metric of choice is optimized at inference time under the distribution from the LLM. We demonstrate across datasets and models how our approach can bring large gains over choosing the most likely target.
@@
Figure 1: An illustration of the metric aware LLM inference. First, an input example is passed to an LLM (forward pass). We next obtain samples from the distribution over targets conditioned on the input $x$ (sampling from $p(y\mid x)$). Finally, the obtained samples are used for finding the target optimizing a metric ${m}$ (metric aware LLM inference); see Table 1 for specific solutions corresponding to example metrics.
2 When (Naïve) LLM Inference Fails
We begin by describing the problem setting. Fix a finite vocabulary $V$ of tokens (e.g., words in English). Let $D$ denote a distribution over inputs $x\in X\subseteq{V}^{*}$ comprising of strings of tokens and targets sequences $y\in Y$. Let $p(y\mid x)$ denote the conditional distribution over targets given an input.
@@ 2 When (Naïve) LLM Inference Fails
A special case of this setting is when $Y\subset\mathbb{R}$ corresponds to numeric targets. Here, we assume that each $y\in Y$ has a unique string representation ${\tt str}(y)\in V^{*}$; for example, the integer $1$ has the string encoding "1". In a slight abuse of notation, we use $p(y\mid x)\stackrel{{\scriptstyle\cdot}}{{=}}p({\tt str}(y)\mid x)$ to denote the conditional probability of output $y$ given input $x$.
@@
A language model (LM) takes a string $x$ as input and predicts an output $\hat{y}\in Y$. Typically, the LM first produces a distribution $\hat{p}(\cdot\mid x)$ over targets, from which a prediction is derived via a suitable inference (or decoding) procedure. Perhaps the most common inference strategy is to choose the mode of $\hat{p}(y\mid x)$:
$\hat{y}(x):=\operatorname*{argmax}_{y\in Y}\hat{p}(y\mid x).$ (1)
In practice, one may approximate the mode by employing greedy decoding or beam search, or sampling multiple candidates with a temperature and picking the among them the one with the highest likelihood score (Naseh et al., 2023).
@@
The quality of an LM’s prediction is measured by some evaluation metric $m(y,\hat{y})$, where we assume that higher values are better. While the exact match (EM), given by $m(y,\hat{y})=\mathds{1}({y=\hat{y}})$, is a commonly used evaluation metric, there are a range of other metrics popularly used to evaluate LMs. These include the (negative) squared error $m(y,\hat{y})=-(y-\hat{y})^{2}$ or absolute error $m(y,\hat{y})=-|y-\hat{y}|$ in the case of numerical targets, or text-based metrics such as $F_{1}$ or BLEU.
A natural goal is to then choose the inference strategy $\hat{y}(x)$ to maximize the metric $m$ of interest, i.e., to maximize the following expected utility:
@@
As we shall see below, for many choices of metric $m(y,\hat{y}(x))$, picking the mode of the predicted distribution (1) can be sub-optimal for (2).
Example 1.
@@
Consider the task of predicting the star rating (on the scale 11115555) associated with a review text. Suppose $m(y,\hat{y})$ is the negative absolute error between true and predicted star ratings. Given the review text ‘‘This keybord is suitable for fast typers’’, suppose the responses and the associated probabilities from an LM are {‘‘1’’: 0.3,  ‘‘2’’: 0.0,  ‘‘3’’: 0.3,  ‘‘4’’: 0.0,  ‘‘5’’: 0.4}. The mode of the predicted probabilities is ‘‘5’’. In contrast, the maximizer of (2) is the median rating ‘‘3’’. We provide real examples for Amazon reviews of learn probability distributions and the label in Figure 2 (Appendix).
Example 2.
Consider a reading comprehension task, where we use the $F_{1}$ score as the evaluation metric $m(y,\hat{y})$, defined by the harmonic mean of $\text{recall}(y,\hat{y})=\frac{|y\cap\hat{y}|}{|y|}$ and $\text{precision}(y,\hat{y})=\frac{|y\cap\hat{y}|}{|\hat{y}|}$. Suppose for the question ‘‘What is the hottest month in the year’’, the responses and associated probability from an LM are {‘‘July’’: 0.25,  ‘‘July 2023’’: 0.23,  ‘‘Month of July’’: 0.24,  ‘‘May’’: 0.28}. The mode of this distribution is ‘‘May’’; whereas the maximizer of (2) is ‘‘July’’.
3 Metric-aware LLM inference
We seek to design decoding strategies that maximize the expected utility in (2).
3.1 Bayes-optimal inference
In an idealized setting, where we have access to the true conditional probabilities $p(\cdot\mid x)$, it is straight-forward to show that the maximizer of (2) is:
$\hat{y}^{*}(x)\in\operatorname*{argmax}_{y^{\prime}\in Y}\,\mathbb{E}_{y\sim p(\cdot\mid x)}\left[m(y,y^{\prime})\right].$ (3)
When $m$ is the EM metric, the optimal inference strategy is $\hat{y}^{*}(x)\in\operatorname*{argmax}_{y\in Y}\,p(y\mid x),$ which is what common approaches such as greedy decoding seek to approximate.
In general, however, the optimal decoding strategy can have a very different form. For example, for evaluation metrics over numerical targets such as the squared error or the absolute error, the optimal inference strategy is to simply take the mean or median of $p(\cdot|x)$ (Bishop, 2006). Table 1 summarizes these optimal rules for different metrics.
3.2 Approximating the optimal solution
In practice, we mimic the Bayes-optimal solution in (3) with two approximations. First, we replace the true conditional distribution ${p}(\cdot\mid x)$ with the LM’s predicted distribution $\hat{p}(\cdot\mid x)$. This is a reasonable approximation when the LM is pre-trained with next-token prediction task and softmax cross-entropy loss: the latter is a strictly proper loss, whose minimizer under an unrestricted hypothesis class is the true conditional distribution $p(y\mid x)$ (Gneiting and Raftery, 2007).
Second, we estimate the expectation in (3) by sampling $K$ outputs from $\hat{p}(\cdot\mid x)$ with a sampling temperature $T$, and then computing:
As noted in Table 1, the above maximization can be computed in closed-form for some common metrics. In applications where we neither have a closed-form solution, nor can enumerate $S$ feasibly, we may solve (4) approximately by maximizing over a fixed subset $C\subset Y$ of candidate outputs:
The choice of $C$ would depend on the task and metric at hand. For example, one simple choice is for $C$ to be the $K$ sampled outputs, i.e., $C=\{y_{1},\ldots,y_{K}\}$. In Appendix A.1, we described how we choose $C$ for text generation tasks.
The quality of our inference strategy depends on the sampling temperature $T$, the number of samples $K$, and the choice of $C$. Note that sampling can be done efficiently by caching the Transformer activations for the input prefix when generating different targets; the Transformer computation needs to be repeated only for the different sampled targets.
3.3 Post-hoc temperature scaling
As mentioned, when sampling from $\hat{p}(\cdot\mid x)$, it often helps to apply a temperature scaling to the LM logits to control the diversity of the sampled outputs. This is particularly important in our procedure where we wish to approximate expectations over $\hat{p}(\cdot|x)$ using a few samples.
In practice, one may sample from $\hat{p}(\cdot\mid x)$ with temperature $T=1$, and apply temperature scaling in a post-hoc manner by employing a weighted version of the objective in (4):
$\hat{y}(x)\in\operatorname*{argmax}_{y^{\prime}\in Y}\,\sum_{i=1}^{K}\left(\hat{p}(y_{i}\mid x)\right)^{\alpha}\cdot m(y_{i},y^{\prime}),$ (6)
where $\alpha$ can be seen as the temperature scaling parameter. The above summation is a (scaled) estimate of $\mathbb{E}_{y\sim\hat{p}(\cdot\mid x)}\left[\hat{p}(y\mid x)^{\alpha}\cdot m(y,y^{\prime})\right]$. For probabilities $\hat{p}(y_{i}\mid x)\propto\exp(f(x,y_{i}))$ defined by logits $f(x,y_{i})$, this is equivalent to computing the expectation under the temperature-scaled distribution $\hat{p}_{\alpha}(y\mid x)\propto\exp((1+\alpha)\cdot f(x,y))$, albeit a normalization factor. We consider an analogous weighting scheme for the plug-in estimators of the closed-form solutions reported in Table 1.
Problem Label space $Y$ Metric Optimal decision rule
Classification $1,\ldots,K$ $\mathds{1}({y=\hat{y}})$ $\hat{y}(x):=\operatorname*{argmax}_{y}[p(y\mid x)]$
Regression $\mathbb{R}$ $-(y-\hat{y})^{2}$ $\hat{y}(x):=\mathbb{E}_{y\sim p(\cdot\mid x)}[y]$
Ordinal regression $1,\ldots,K$ $-|y-\hat{y}|$ $\hat{y}(x):=\text{median}[p(\cdot\mid x)]$
Question answering $V^{*}$ $F_{1}(y,\hat{y})$ $\hat{y}(x):=\operatorname*{argmax}_{y^{\prime}}\,\mathbb{E}_{y\sim p(\cdot\mid x)}\left[m(y,y^{\prime})\right]$
Bi-partite ranking ±$\pm 1$ AUC $\hat{y}(x):=p(y=+1|x)$
Multi-partite ranking $1,\ldots,K$ $\text{AUC}^{\text{multi}}$  with costs $c_{y,y^{\prime}}=|y-y^{\prime}|$ $\hat{y}(x):=\mathbb{E}_{y\sim p(\cdot\mid x)}[y]$
Table 1: Optimal inference for different evaluation metrics. For the first three, we have a closed-form solution. For $F_{1}$, we approximate the optimal solution by optimizing over a subset of candidates $C\subset Y$3.2). See (§3.5) for definitions of AUC. For the problems in rows 1–4, the LLM predicts an output in label space $Y$; for the ranking problems in rows 5–6, the LLM predicts a real-valued score, which is then used to rank the samples.
model size greedy decode MALI
argmax mean $*{}^{*}$
STSB XXS 1.078 1.448 1.007 1.028
S 0.685 1.019 0.641 0.649
L 0.628 0.989 0.580 0.610
argmax median $*{}^{*}$
Amazon reviews XXS 0.495 0.826 0.493 0.474
S 0.301 0.444 0.299 0.285
L 0.294 0.541 0.298 0.291
argmax $F_{1}$ ${F_{1}}^{*}$
Trivia-QA XXS 0.314 0.178 0.307 0.302
S 0.620 0.636 0.650 0.678
L 0.886 0.887 0.880 0.886
Table 2: Root mean squared error (RMSE) on STSB dataset (lower is better), Mean absolute error (MAE) on Amazon reviews dataset (lower is better), and $F_{1}$ metrics on Trivia-QA dataset (higher is better) from PaLM-2 models of varying size. We sample $16$ targets for the argmax and MALI methods. All methods rely on $T=1$ temperature for sampling, and the weighted approaches correspond to an effective temperature $T=\frac{1}{4}$. See Tables 5, 6, 7 in the Appendix for results over more temperatures and evaluation metrics (cost weighted multi-class AUC and Pearson correlation).
3.4 Efficiency
Note that sampling can be done efficiently by caching the Transformer activations for the input prefix when generating different targets. The Transformer computation needs to be repeatedly done only for the different sampled targets.
In practice, when the prefix is long compared to the generated targets (e.g. for scoring tasks, where the target score can be just a few tokens length, whereas the prefix can be long as it contains the input text), a forward pass for the prefix tends to take most of the compute time.
3.5 Ranking metrics
Apart from the settings where evaluation metrics are decomposable, i.e., evaluation metrics are a result of aggregation of per-example results from evaluation, there are also metrics which are non-decomposable. An example for such a metric are ranking metrics, where the goal is to correctly order examples, as opposed to predict correct scores or classes. Ranking metrics are appropriate for many important tasks, including ads click-through rate prediction or recommender systems. Below we review different types of ranking metrics and the corresponding optimal decision rules.
Bi-partite ranking
In bi-partite ranking, the goal is to order examples from two classes (and so the label set $Y={\pm 1}$) in such a way that examples from the positive class are above the examples from the negative class. In this case The LLM predicts real valued scores, and so the target space need not be the same as the label space. One common metric used for evaluating bi-partite ranking is the area under the ROC curve (AUC), which is proportional to the number of wrongly ordered pairs of examples in the ranking (Cortes and Mohri, 2003):
$\text{AUC}(\hat{y})\,=\,1-\mathbb{E}\Big{[}\mathds{1}(\hat{y}(x)<\hat{y}(x^{\prime}))\,+\,\frac{1}{2}\cdot\mathds{1}(\hat{y}(x)=\hat{y}(x^{\prime})\,\Big{|}\,y=+1,y^{\prime}=-1\Big{]}$
It can be shown that the optimal decision rule for AUC takes the form of (any strictly monotone transformation of) $\hat{y}^{*}(x)=p(y=+1|x)$ (Clémençon et al., 2008; Uematsu and Lee, 2015).
Multi-partite ranking
Multi-partite ranking is an extension to the bi-partite ranking, where the ranking is over examples from multiple categories. Several multi-category extensions of binary AUC have been considered, for example by counting wrongly ordered pairs of examples in the ranking according to the category indices and weighing them according to predetermined costs (Uematsu and Lee, 2015):
$\text{AUC}^{\text{multi}}(\hat{y})\,=\,1-\mathbb{E}\Big{[}c_{y,y^{\prime}}\cdot\Big{(}\mathds{1}(\hat{y}(x)<\hat{y}(x^{\prime}))\,+\,\frac{1}{2}\cdot\mathds{1}(\hat{y}(x)=\hat{y}(x^{\prime})\Big{)}\,\Big{|}\,y>y^{\prime}\Big{]}$
Uematsu and Lee show that when the costs are equal to the difference between category indices, i.e., $c_{y,y^{\prime}}=|y-y^{\prime}|$, an optimal solution is to use mean distribution over scores: $\hat{y}^{*}(x)=\mathbb{E}_{y\sim p(\cdot|x)}\left[y\right]$. We report results on the cost weighted multi-class AUC in Table 6 in Appendix.
4 Experiments and discussion
We compare our metric-aware inference approach against standard inference strategies on NLP tasks with different evaluation metrics.
Datasets. We summarize the datasets below, and list the prompts used in Table 4 (Appendix).
  1. (i)
    STSB: The Semantic Textual Similarity Benchmark (Cer et al., 2017) comprises of sentence pairs human-annotated with a similarity score from 0 to 5. As it is a regression task, we consider evaluation with root mean squared error.
  2. (ii)
    US Amazon reviews: We evaluate on a sample of $1500$ examples from the US Amazon reviews dataset. From the review text we aim to predict the 5-star rating corresponding to a product review (Ni et al., 2019). Since the task is in the form of ordinal regression, we use mean absolute error as the metric (Fathony et al., 2017).
  3. (iii)
    TriviaQA is a reading comprehension dataset comprising (question, answer, reference text) triplets, where a question is answered against a refence text (Joshi et al., 2017). The evaluation is the $F_{1}$ score (see Example 2).
We use a sample of $1500$ examples for evaluation.
Models. We consider two instruction-tuned model families: FLAN-T5 (Chung et al., 2022) and PaLM-2 (Google and et al., 2023). We report results across different model sizes and temperatures. Unless otherwise stated, we fix the number of samples to $K=16$, and the top-$k$ parameter in decoding to $40$ (Fan et al., 2018)
Methods. We evaluate the following methods (i) greedy decoding, (ii) sampling $K$ candidates and picking the one with the maximum likelihood (argmax), (iii) the proposed MALI approach on the same $K$ samples, and (iv) the temperature scaled variant of MALI in §3.3 (denoted by a ‘*’). For (iv), we choose $\alpha$ so that the effective temperature is $\frac{1}{4}$. For the tasks involving the squared and absolute error metrics, we use the closed-form decision rules in Table 1. For the $F_{1}$ metric, we solve (5) over a candidate set $C$, which we choose to contain the $K$ samples and additional targets derived from them (details in Appendix A.2).
Metric-aware inference helps. In Table 2, we report results across datasets and model sizes. We notice that MALI improves over baselines across all model sizes on STSB and Amazon reviews (for model size S, we see that median performs very similarly to the most likely generated sample). On Trivia QA, we find that on XXS and S, optimizing for $F_{1}$ helps, and for L, simply using the most likely generated sample works most effectively.
We see diminishing improvements with larger models for Trivia QA and Amazon datasets. This coincides with a lowering entropy in predictions with increasing model size for Trivia QA and Amazon datasets, whereas STSB does not exhibit this phenomenon as strongly: see Table 8 (Appendix) for empirical entropy estimates.
We also evaluate MALI with varying number of samples $K$. As seen in Table 9 (Appendix), on STSB with temperature $\frac{1}{4}$, even with as few as two samples, MALI starts to show improvements over greedy decoding.
Sampling versus scoring. So far, when estimating the prediction maximizing the (2), we have used sampling from the LM distribution (see §3.2). Alternatively, if the targets are from a narrow interval (e.g., on the STSB dataset, the values are in the interval $[0,5]$), one can find the probabilities for targets at fixed intervals (e.g. $0,0.5,1.0,\ldots,5.0$), and thus find the solutions from Table 1 and more generally from (3).
In Table 3, we report results from FLAN-T5 on the STSB dataset for both scoring and sampling based estimates for MALI, where the scoring based estimator is based on $11$ equally spaced targets. We find that both sampling and scoring lead to MALI improving over choosing the most likely prediction. Further, we note that sampling is a more effective strategy than scoring the equally spaced targets.
model greedy scoring sampling
FLAN-T5 S 4.419 2.407 2.275
FLAN-T5 L 0.455 0.410 0.373
FLAN-T5 XL 0.508 0.549 0.457
Table 3: Squared error (SE) on STSB from FLAN-T5 models of varying size. Temperature $0.5$ was used for generating targets for the sampling approach.
Conclusions In this work we show how utilizing the output distribution modeled by LLMs in the form of our MALI methods can bring improvements across NLP tasks. Our work also points at importance of further understanding and improving of the quality of model output distribution and calibration, which we found crucial for the MALI methods to work well.
5 Limitations
There are multiple limitations of our work. First, we evaluate our proposed methods on multiple text datasets with numerical and text targets, however, many more types of outputs can be considered, including time series targets, long text targets etc. Next, it would be interesting to more systematically analyze how to solve the objective from ((6)) over many samples for text outputs for metrics like $F_{1}$ or BLEU, e.g. by means of dynamic programming. We also note that the datasets considered in this work are restricted to English. It would be interesting to expand the explorations to datasets in other languages.
6 Ethics Statement
All datasets used in this work are publicly available. No additional user data was collected or released as part of this work. All models used are publicly available and already pretrained, and no finetuning was conducted for any experiments. Instead, all experiments relied on running inference experiments with the models over several thousands of examples. Thus, the CO-2 footprint of this paper is minimal. We do not foresee any significant risks associated with this paper other than improving performance on tasks which are harmful.
7 Acknowledgements
We are thankful to Ziwei Ji for the helpful feedback.
Appendix A Appendix
A.1 Candidate set construction
One simple choice for the candidate set $C$ could be take the $K$ sampled outputs, i.e., $C=\{y_{1},\ldots,y_{K}\}$. One may additionally include in this set transformations on each $y_{i}$ $y_{i}$ or new candidates formed from combining two or more of the samples.
For example, in applications such as reading comprehension or question-answering, where the output is list of keywords that constitute an answer to a question, one often uses a metric such as F1 that balances between precision and recall. In such cases, after sampling, one may additionally include samples formed by concatenating pairs of sampled outputs, i.e., ${\tt concat}(y_{i},{\tt delim},y_{j}),\forall i\neq j$. These concatenated answers have the effect of increasing recall, at the cost of lower precision. We follow that procedure for the Trivia-QA experiments in §4.
A.2 Additional details
Dataset Prompt
STSB
What is the sentence similarity between the following two sentences measured on a scale of 0 to 5: {Sentence #1}, {Sentence #2}. The similarity measured on a scale of 0 to 5 with 0 being unrelated and 5 being related is equal to
Amazon reviews
What is the rating corresponding to the following review in the scale of 1 to 5, where 1 means negative, and 5 means positive? Only give a number from 1 to 5 with no text. Review: {Review} Rating:
Trivia-QA
Answer the following question without any additional text. Question: {Question}. Answer:
A.3 Additional experiments
In Table 8 we report empirical entropy estimates as measured based on the $16$ samples generated from the model. We find that entropy decreases as model size increases. We observe a particularly sharp decrease in entropy for the Amazon reviews and Trivia-QA datasets, where for larger model sizes we don’t find improvements from MALI approaches.
In Table 5 we report RMSE on STSB dataset, MAE on Amazon reviews dataset, and $F_{1}$ metrics on Trivia-QA dataset from PaLM-2 models of varying size across multiple temperature values. We find improvements over baselines on STSB and Amazon reviews datasets for most temperatures. For Trivia-QA, we find improvements for XXS and S models for some temperatures, and for L, we don’t find a difference from our methods due to low entropy in the responses (see Table 8). In Table 7 we additionally report Pearson correlation metrics on STSB, confirming the results of MALI improving over autoregressive inference. Lastly, in Table 6 we report cost weighted multi-class AUC with costs corresponding to the difference between the annotated labels: $|y_{1}-y_{2}|$. We find on both STSB and Amazon reviews datasets that the optimal decision rule (mean over the distribution) improves over the baselines.
In Table 9, we report the impact of the number of samples on the results. We note that there is an improvement in the results with the increase in the number of samples, however beyond 8 samples there is a diminishing improvement in practice.
In Figure 2 we report examples from the Amazon dataset and the corresponding: human annotations and samples from the model. Notice how samples cover significant proportions of the ratings. We find that the samples end up in the vicinity of the human annotation, and thus in many cases taking a mean over samples helps improve the prediction over the mode.

### v3

# Regression-aware Inference with LLMs

@@ Abstract
Large language models (LLMs) have shown strong results on a range of applications, including regression and scoring tasks. Typically, one obtains outputs from an LLM via autoregressive sampling from the model’s output distribution. We show that this inference strategy can be sub-optimal for common regression and scoring evaluation metrics. As a remedy, we build on prior work on Minimum Bayes Risk decoding, and propose alternate inference strategies that estimate the Bayes-optimal solution for regression and scoring metrics in closed-form from sampled responses. We show that our proposal significantly improves over baselines across datasets and models.
@@ 1 Introduction
Large language models (LLMs) are currently the most capable models across many NLP tasks (OpenAI et al., 2023; Anil et al., 2023; Touvron et al., 2023; Gemini Team et al., 2024). Owing to their remarkable few- and zero-shot abilities (Wei et al., 2022; Kojima et al., 2023), pre-trained LLMs are often applied without any additional training on domain-specific datasets: instead, one may query the LLM with a suitably crafted input prompt.
More recently, LLMs have been successfuly applied to regression and scoring tasks. For example, Gruver et al. (2023) explored zero-shot learning for time series prediction; Vacareanu et al. (2024) showed how LLMs are remarkably strong at in-context learning for regression tasks; Liu and Low (2023); Yang et al. (2023) considered autoregressive fine-tuning over numerical targets applied to arithmetic tasks; and Qin et al. (2023) applied LLMs for listwise ranking.
The quality of an LLM is often assessed using an application-specific evaluation metric. One popular metric is the exact match (EM), which penalises any response not exactly equal to the one in the dataset annotation. This is an analogue of the conventional classification accuracy. While EM is an intuitive metric, there are many applications where it is not suitable. This includes tasks such relevance scoring (Cer et al., 2017) and sentiment analysis (Fathony et al., 2017), where the outputs are numerical or ordinal categories. In these cases, one instead prefers metrics such as the squared error, absolute error or ranking scores that take the outputs’ ordinal nature into account.
@@
Despite the wide variety of evaluation metrics, LLM inference is typically performed in the same manner for every task: namely, one performs auto-regressive sampling from the LLM’s underlying distribution (see §2). While intuitive, such inference does not explicitly consider the downstream evaluation metric of interest. This raises a natural question: is there value in adapting the inference procedure to the evaluation metric at hand for regression and scoring tasks?
A prominent line of work takes a decision-theoretic approach to the above problem. Dubbed as Minimum Bayes Risk (MBR) decoding, this approach seeks to optimize at inference time the metric of choice under the model’s distribution (Bickel and Doksum, 1977; Kumar and Byrne, 2004; Eikema and Aziz, 2020; Bertsch et al., 2023). Much of the work on MBR is focused on evaluation metrics for machine translation and text generation tasks, such as the BLEU score (Papineni et al., 2002). Of particular interest in this literature are self-consistency based decoding strategies that take a (weighted) majority vote of sampled responses (Wang et al., 2023a), which have shown to provide quality gains in arithmetic and reasoning problems.
In this paper, we build on the existing literature on MBR to design metric-aware inference strategies for general regression and scoring tasks. We first observe that choosing the most likely target for an input corresponds to inherently optimizing for the EM metric, and is consequently not optimal when EM is not the metric of choice. As a remedy, we propose estimating the Bayes-optimal output for a metric under the model’s distribution (see Figure 1 for an illustration of our method); we show that this admits a closed-form solution for common regression and ranking metrics, and only requires estimating a simple statistic from the sampled responses. In contrast, prior MBR methods for translation and summarization often require heuristically solving an intractable maximization problem (Ehling et al., 2007; Bertsch et al., 2023). We show across datasets and models how our approach yields gains over choosing the most likely target, and over self-consistency based approaches.
@@
Figure 1: Illustration of metric-aware LLM inference for regression and scoring tasks. An input $x$ is passed to the LLM, and samples are drawn from the distribution over targets $y$ conditioned on $x$. These are then used to find the target optimizing a metric ${m}$ through a closed-form decision rule $\Phi$ (e.g., mean or median); Table 1 presents specific solutions across metrics.
Algorithm 1 RAIL: Regression-aware Inference with LLMs
1:input: Model $M$, #samples $K$, sampling temperature $T$, effective temperature $T^{\prime}$, decision rule $\Phi$
2:for $i=1$ to $K$ do
3:     $\hat{y}_{i},\hat{p}_{i}=M.\text{\tt generate}(T)$$\triangleright$ $M$.generate returns the sampled target and its probability.
4:end for
5:$\alpha=\frac{T}{T^{\prime}}-1$$\triangleright$ Compute post-hoc temperature scaling so that the effective temperature used is $T^{\prime}$
6:return $\Phi(\hat{y}_{1},\ldots,\hat{y}_{K},\hat{p}_{1},\ldots,\hat{p}_{K},\alpha)$
2 When (naïve) LLM inference fails on regression tasks
@@
We begin with the problem setting. For a finite vocabulary $V$ of tokens (e.g., words in English), let $D$ denote a distribution over inputs $x\in X\subseteq{V}^{*}$ comprising of strings of tokens, and targets $y\in Y$. Let $p(y\,|\,x)$ denote the conditional distribution over targets given an input. We consider a special case of this setting where $Y\subset\mathbb{R}$ corresponds to numeric targets. Here, we assume that each $y\in Y$ has a unique string representation ${\tt str}(y)\in V^{*}$; for example, the integer $1$ has the string encoding "1".
A language model (LM) takes a string $x$ as input and predicts an output $\hat{y}\in V^{*}$. Typically, the LM first produces a distribution $\hat{p}(\cdot\,|\,x)$ over targets. In a slight abuse of notation, we use $\hat{p}(y\,|\,x)\stackrel{{\scriptstyle\cdot}}{{=}}\hat{p}({\tt str}(y)\,|\,x)$ to denote the conditional probability of a numerical output $y$ given input $x$. Note that even for problems where numerical targets are expected, an LM may return a non-zero probability to non-numerical targets.
@@
A prediction from an LM is typically derived via a suitable inference (or decoding) procedure. Perhaps the most common inference strategy is to choose the mode of $\hat{p}(\cdot\,|\,x)$:
$\hat{y}(x):=\operatorname*{argmax}_{y\in V^{*}}\hat{p}(y\,|\,x).$ (1)
In practice, one may approximate the mode via greedy decoding or beam search, or sampling multiple candidates and picking the among them the one with the highest likelihood score (Naseh et al., 2023). In principle, the extracted target may not be numerical. In such cases, a possible strategy is to resort to predicting a default numerical value such as $0.0$. In practice, we find the targets from high-quality LLMs tend to be numerical even under zero-shot settings, and so converting most likely targets from $V^{*}$ to $Y$ is usually possible.
@@
The quality of an LM’s prediction is measured by some evaluation metric $m(y,\hat{y})$, where we assume that higher values are better. While the exact match (EM), given by $m(y,\hat{y})=\mathds{1}({y=\hat{y}})$, is a commonly used evaluation metric, there are a range of other metrics popularly used to evaluate LMs. These include the (negative) squared error $m(y,\hat{y})=-(y-\hat{y})^{2}$ or absolute error $m(y,\hat{y})=-|y-\hat{y}|$ for regression tasks. A natural goal is to then choose the inference strategy $\hat{y}(x)$ to maximize the metric $m$ of interest, i.e., to maximize the expected utility:
@@
For many choices of metric $m(y,\hat{y}(x))$, picking the mode of the predicted distribution (1) can be sub-optimal for (2). As an example, consider predicting the star rating (on the scale 11115555) associated with a review text. Suppose $m(y,\hat{y})$ is the negative absolute error between the true and predicted ratings. Given the review text ‘‘This keybord is suitable for fast typers’’, suppose the LM responses and the associated probabilities are {‘‘1’’: 0.3,  ‘‘2’’: 0.0,  ‘‘3’’: 0.3,  ‘‘4’’: 0.0,  ‘‘5’’: 0.4}. The mode of the predicted probabilities is ‘‘5’’. In contrast, the maximizer of (2) is the median ‘‘3’’.